How many litres of 12% sodium hypochlorite solution are needed to make 100 litres of 2% sodium hypochlorite solution?
Answer by mananth(16946)

0.12x = (100+x)0.02
0.12x=1+0.02x
multiply by 100
12x=100+2
12x=102
x=8.5
8.5 liters